jonacriezl:

Show that if the volume of a balloon is decreasing at a rate proportional to its surface area, the radius of the balloon is shrinking at a constant rate.

jhonyy9:

- hope you know that a balloon has a geometric form of a sphere - so in this case you need to know formula for the volume of a sphere and formula for surface area of a sphere - using these two formulas you need prove this rate proportional

Vocaloid:

My calculus is a bit rusty so bear with me Since we’re dealing with rates of change, we’ll be using derivatives dV/dt gives the volume change wrt t Since the rate of decrease is proportional to surface area, dV/dt = k(SA) Surface area of a sphere is 4(pi)(r)^2 Therefore dV/dt = k * 4pi(r^2) Find dr/dt through implicit differentiation wrt t and you’ll notice that dr/dt ends up being a constant if you do it correctly

Timmyspu:

Then what variable would you be solving for?

Vocaloid:

It’s not asking for a variable, it’s asking to prove the radius rate of change is a constant

jhonyy9:

@Vocaloid yes i agree it but i start - begin - it always that who asked dont learned again about derivare - how calcule the derivative

jhonyy9:

so in this situation how you solve it ?

Vocaloid:

How would you approach this without derivatives ? (Asking our of curiosity)

Timmyspu:

You would have to solve using a equation and @jhonyy9 knows how to do that equation.

jhonyy9:

- this is true that in my birth country i ve learned about derivatives in my 3rd class of my high school but here in Hungary dont learn about derivare just in university or faculty - college

Timmyspu:

Ok so the first thing is you need to have a example equation to reference off of because that will help you a lot.

Timmyspu:

Also what @Vocaloid said above about having a constant if you dont have a constant then every time you will get the answer wrong.

Vocaloid:

@jhonyy9 would you mind sharing the non calculus solution sometime? (You can pm me if you don’t want to give away the answer)

jhonyy9:

V = (4pi*r^3)/3 A = 4pi*r^2 \[\frac{ V }{ A } = \frac{ \frac{ 4\pi*r^3 }{ 3 } }{ 4\pi*r^2 }\] \[\frac{ V }{ A } = \frac{ r }{ 3}\] @Vocaloid

Vocaloid:

That just shows that volume and surface area are proportional for constant r. I don’t see how that proves that r changes at a constant rate when the volume rate of change is proportional to SA.

jhonyy9:

maybe this with \[\frac{ \delta V }{ \delta A } = \frac{ \delta r }{ 3 }\] \[\frac{ V_2 - V_1 }{ A_2 - A_1 } = \frac{ (r_2 -r_1) }{ 3 }\]

jhonyy9:

so result that the constant is 1/3

jhonyy9:

@Vocaloid really i agree what you ve said above with derivative but how you solve this problem when the student who asked dont learn again about derivare ?

Vocaloid:

Ok thank you, I was genuinely just curious about your method

jhonyy9:

@Timmyspu

jhonyy9:

np yw like a first idea i ve thought it like you but i always i think in this second way what is in case then who asked dont learn again about calculus - in this case how maybe possibile solve the posted problem ??? thank you all your collaboration

Timmyspu:

No problem!

SemiDefinite:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @jhonyy9 maybe this with \[\frac{ \delta V }{ \delta A } = \frac{ \delta r }{ 3 }\] \[\frac{ V_2 - V_1 }{ A_2 - A_1 } = \frac{ (r_2 -r_1) }{ 3 }\] \(\color{#0cbb34}{\text{End of Quote}}\) this is nonsense in so many ways

SemiDefinite:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @jhonyy9 so result that the constant is 1/3 \(\color{#0cbb34}{\text{End of Quote}}\) sheesh, so it's always 1/3, fixed in time forever?!

jhonyy9:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @jhonyy9 V = (4pi*r^3)/3 A = 4pi*r^2 \[\frac{ V }{ A } = \frac{ \frac{ 4\pi*r^3 }{ 3 } }{ 4\pi*r^2 }\] \[\frac{ V }{ A } = \frac{ r }{ 3}\] @Vocaloid \(\color{#0cbb34}{\text{End of Quote}}\) look here pls

jhonyy9:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @jhonyy9 maybe this with \[\frac{ \delta V }{ \delta A } = \frac{ \delta r }{ 3 }\] \[\frac{ V_2 - V_1 }{ A_2 - A_1 } = \frac{ (r_2 -r_1) }{ 3 }\] \(\color{#0cbb34}{\text{End of Quote}}\) and continue here

jhonyy9:

and from this result this 1/3 constant

SemiDefinite:

you were supposed to show that the answer is k, the constant. not 1/3. a very specific answer, don't you think?

jhonyy9:

in this case of this above posted problem the k constant will be always k= 1/3 check it please

SemiDefinite:

\(V = {4 \over 3} \pi r^3 \qquad S = 4 \pi r^2\) \(dV = 4 \pi r^2 ~ dr\qquad dS = 8 \pi r ~ dr\) \(\frac{dV}{dt} = - k S\) \(\implies \frac{4 \pi r^2 ~ dr}{dt} = - k (4 \pi r^2)\) \(\implies \frac{ dr}{dt} = - k\) That is what you are supposed to demonstrate. What you are posting here is pure nonsense, sadly.

SemiDefinite:

\( d\left( \dfrac{A}{B} \right) \ne \dfrac{dA}{dB} \) think about it.